Advanced Data Analysis II

36-402 Spring 1996

Hints on the final

Many of you presented very good preliminary analyses of the data. I have tried to put some specific comments on each page, but there are also two generic comments.

• You can probably do better for some models by breaking the data into different soil types. There is no good reason to believe that the same model will work for all types of soil, so you should at least allow for the possibility of different models, either by splitting the data, or building models with soil interactions.

• Be very careful you don’t use something as a predictor which can’t be known until after the data are collected. (To be blunt, don’t use one of the response variables as a predictor for another model).